OpenStudy (anonymous):

Solve. 7x^2 - 63 = 0 x = 7 and x = 9 x = 9 and x = 3 x = 3 and x = -3 x = 3 and x= 7

OpenStudy (mertsj):

3 and -3

hero (hero):

7x^2 - 63 = 0 7x^2 = 63 x^2 = 63/7 x^2 = 9 x = +3 x = -3

OpenStudy (pokemon23):

"hero link down below" http://openstudy.com/study#/updates/4f14ec5ee4b0b9109c921eb8